OpenStudy (anonymous):

a water tank that hold 60 L of water can be emptied in 24 min. how long will it take to empty a water tank that holds 280 L of water?

OpenStudy (anonymous):

Im going to say subtract 60 from 280 and then divide by time.

OpenStudy (anonymous):

24/60=x/280 60x=6720 x=112

OpenStudy (primeralph):